OpenStudy (anonymous):

how do I find the domain and intercepts of f(x)= ln(x+2)

OpenStudy (therealmeeeee):

ok x - 1 > 0 x > 1 y > 2 x-int: ln (x - 1) + 2 = 0 ln (x - 1) = -2 x - 1 = e^(-2) x = 1 + e^(-2) no y-int.

OpenStudy (therealmeeeee):

did this help???????????

OpenStudy (therealmeeeee):

omd dude this help